the perimeter is 46cm and length of one side is 8cm find the length of the diagonal.

Answer by oberobic(2304)

Assuming this is a rectangle,
P = 46
P = 2(L+W)
L+W = 23
L = 8
W = 23-8 = 15
.
Assuming this is a rectangle, we can use the Pythagorean Theorem to find the length of the diagonal.
.
c^2 = 8^2 + 15^2
c^2 = 64 + 225
c^2 = 289
c = sqrt(289)
c = 17
.
The sqrt(289) = + or -17, but we can ignore the -17 because a negative length for the diagonal is nonsense.
.
Answer: The diagonal is 17 cm.